Edgewater Enterprises manufactures two products. Information follows:     

Product A Product B
Sales price $ 13.70 $ 17.50
Variable cost per unit $ 6.70 $ 7.20
Product mix 40% 60%


Calculate the break-even point if Edgewater’s total fixed costs are $235,000. (Round your intermediate calculations to 2 decimal places and final answer to the nearest whole number.)
Answer #1
A B
Contribution margin=(Sales price-Variable cost) (13.7-6.7)=$7 (17.5-7.2)=$10.3

Weighted average Contribution margin=Respective Contribution margin*Respective Sales Mix

=(7*0.4)+(10.3*0.6)=$8.98

Hence breakeven point=Fixed cost/Weighted average Contribution margin

=(235000/8.98)

=26,169 units(Approx).